An express train and a local train leave Xanadu at 3p.m. and head for Nirvana 50 miles away. The express travels twice as fast as the local and arrives an hour ahead of the local. What is the speed of each train?
Answer by mananth(16949)

speed of local train =x mph
speed of express = 2x
..
d= 50
time local - time express =1
50/x - 50/2x= 1
(100-50) /2x=1
50=2x
x= 25 mph local train
express train speed = 50 mph
...
m.ananth@hotmail.ca